# Bounds and Constructions of High-Memory Spatially-Coupled Codes

### v1

@@ III-B Theoretical Analysis of Algorithms
Algorithm 1 The Moser-Tardos Algorithm [25]
0:  Sample space: $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$.Independent r.v. taking values in $\mathbf{a}$: $\chi=\{X_{1},...,X_{\gamma\kappa}\}$.Set of (ordered) events: =$\mathcal{H}=\{H_{1},...,H_{k}\}.$probability distribution: $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$.
@@
1:  Sample the variables $X_{i},i\in[\gamma\kappa]$, and let $\alpha$ be the resulting assignment.
@@
[…]
@@ IV Characterization of the Solution Space
The distribution of $\Omega$ conditioned on avoiding $\mathcal{A}$ is called the LLL-distribution. The distribution at the M-T algorithm termination is called the MT-distribution.
@@
If the conditions in asymmetric LLL[27, Theorem 1.1] are satisfied, then the LLL-distribution $D_{LLL}$ is well-defined. For any event $E$ determined by the set of random variables $\chi$, the probability $P_{LLL}[E]$ under the LLL-distribution $D_{LLL}$ satisfies:
@@
Specifically, for symmetric LLL, there is a more intuitive result.
Corollary 3
If the conditions in symmetric LLL[27, Theorem 1.5] are satisfied, then the LLL-distribution $D_{LLL}$ is well-defined. For any event $E$ determined by the set of random variables $\chi$, the probability $P_{LLL}[E]$ under the LLL-distribution $D_{LLL}$ satisfies:
@@
[26] also proves that the same conclusion holds for the MT-distribution, indicating that the MT-distribution can effectively approximate the LLL-distribution.
Lemma 6
If the conditions in asymmetric LLL[27, Theorem 1.1] are satisfied, then the MT-distribution $D_{MT}$ is well-defined. For any event $E$ determined by the set of random variables $\chi$, the probability $P_{MT}[E]$ that $E$ was true at least once during the execution of the MT-algorithm on the events in $\mathcal{A}$, is at most $P_{\Omega}[E]\cdot\prod_{B\in\mathcal{N}(E)}(1-x_{B})^{-1}$. In particular, the probability of $E$ being true in the output distribution of MT obeys this upper-bound.
$P_{MT}[E]\leq P_{\Omega}[E]\cdot\prod_{B\in\mathcal{N}(E)}(1-x_{B})^{-1}$ (26)
The two lemmas above indicate that when an event $E$ is associated with relatively few bad events $A\in\mathcal{A}$, its probability of occurring does not increase significantly.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. Suppose =$\mathcal{H}=\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{f}}\}$ are the probabilities that they are active after the random edge-spreading and lifting process. If =$\mathcal{H}^{\prime}=\{H_{1},...,H_{f}\}\subseteq\mathcal{H}=\{H_{1},...,H_{k}\}$ with dependency graph $G^{\prime}=G[\mathcal{H}^{\prime}]$ satisfies the conditions in asymmetric LLL[27, Theorem 1.1], then for any event $E$ determined by the set of random variables $\chi$, the probability $P_{LLL}[E]$ under the LLL-distribution $D_{LLL}$ satisfies:
@@
And the probability of $E$ being true in the output distribution of MT also obeys this upper-bound.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges. If $\mathcal{H}^{\prime}=\mathcal{C}_{4}$ satisfies the conditions in symmetric LLL[27, Theorem 1.5], then $P_{LLL}[c_{2k}]\leq P_{\Omega}[c_{2k}]\cdot(1+\frac{1}{\Delta})^{2kW_{\mathcal{H}^{\prime}}}$. Specifically, $P_{LLL}[c_{6}]\leq P_{\Omega}[c_{6}]\cdot e^{8/3}$
@@
According to Corollary 3, we have
@@
$\leq P_{\Omega}[c_{6}]\cdot e^{8/3}$ (32)
V CONCLUSION
This paper primarily employs probabilistic combinatorial methods to study the construction of QC-SC-LDPC codes. To achieve strong performance in the error-floor region, it is necessary to eliminate certain harmful combinatorial structures that degrade decoding. By formulating this problem within the framework of the Lovász Local Lemma (LLL), we derive upper bounds on both memory and lifting degree, and propose a construction algorithm with theoretically predictable performance. In addition, we investigate the mutual influence between the elimination of different harmful structures, as previously observed in experiments, and provide explicit upper bounds on these effects.
Acknowledgment
This work is partially supported by the National Key R&D Program of China, (2023YFA1009600).

### v2

@@ III-B Theoretical Analysis of Algorithms
Algorithm 1 The Moser-Tardos (MT) Algorithm [26]
0: Sample space: $\mathbf{a}=(a_{0},a_{1},\ldots,a_{m_{t}})$.Independent r.v. taking values in $\mathbf{a}$: $\chi=\{X_{1},...,X_{\gamma\kappa}\}$.Set of (ordered) events: $\mathcal{H}=\{H_{1},...,H_{k}\}.$Probability distribution: $\mathbf{p}=(p_{0},p_{1},\ldots,p_{m_{t}})$.
@@
1: Sample the variables $X_{i},~i\in[\gamma\kappa]$, and let $\alpha$ be the resulting assignment.
@@
2:while there is a least indexed bad event $H_{l}$, such that sc($H_{j}$) $\cap$ sc($H_{l}$) $\neq\emptyset$, occurring under the current assignment, do
@@ IV Characterization of the Solution Space
The distribution of $\Omega$ conditioned on avoiding $\mathcal{A}$ is called the LLL-distribution. The distribution at the MT algorithm termination is called the MT-distribution.
@@
([25]) If the conditions in asymmetric LLL[28, Theorem 1.1] are satisfied, then the LLL-distribution $D_{LLL}$ is well-defined. For any event $E$ determined by the set of random variables $\chi$, the probability $P_{LLL}[E]$ under the LLL-distribution $D_{LLL}$ satisfies:
@@
Specifically, for symmetric LLL[28, Theorem 1.5], there is a more intuitive result.
@@
Moreover, it was shown in [27] that the MT distribution $D_{MT}$ also satisfies the above bounds. The lemma above indicates that when an event $E$ is associated with relatively few bad events $A\in\mathcal{A}$, its probability of occurring does not increase significantly.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges, under the coupling pattern $\mathbf{a}=(a_{0},a_{1},...,a_{m_{t}})$ and corresponding probability distribution $\mathbf{p}=(p_{0},p_{1},...,p_{m_{t}})$. Suppose $\mathcal{H}=\{H_{1},...,H_{k}\}$ are a series of avoidable harmful structures in the base graph with dependency graph $G$, and $\{P_{H_{1}},...,P_{H_{k}}\}$ are the probabilities that they are active after the random edge-spreading and lifting process. If $\mathcal{H}^{\prime}=\{H_{1},...,H_{f}\}\subseteq\mathcal{H}=\{H_{1},...,H_{k}\}$ with dependency graph $G^{\prime}=G[\mathcal{H}^{\prime}]$ satisfies the conditions in asymmetric LLL[28, Theorem 1.1], then for any event $E$ determined by the set of random variables $\chi$, the probability $P_{LLL}[E]$ under the LLL-distribution $D_{LLL}$ satisfies:
@@
And the probability of $E$ being true in the output distribution of Algorithm 1 also obeys this upper bound.
@@
For any QC-SC-LDPC code constructed with respect to a $\gamma\times\kappa$ fully-connected block code base graph, with no parallel edges. If $\mathcal{H}^{\prime}=\mathcal{C}_{4}$ satisfies the conditions in symmetric LLL[28, Theorem 1.5], then $P_{LLL}[c_{2k}]\leq P_{\Omega}[c_{2k}]\cdot(1+\frac{1}{\Delta})^{2kW_{\mathcal{H}^{\prime}}}$. Specifically, $P_{LLL}[c_{6}]\leq P_{\Omega}[c_{6}]\cdot e^{8/3}$.
@@
According to Lemma 5, we have
@@
$\leq P_{\Omega}[c_{6}]\cdot e^{8/3}~(\gamma,\kappa\geq 3)$ (31)